OpenStudy (anonymous):

Match each expression with its solution. Match Term Definition fifty point five divided by zero point five A) ten point one twenty−six point nine minus eleven point six B) one hundred one one hundred seventy times zero point nine C) one hundred fifty−three three point nine plus six point two D) fifteen point three

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

50.5 / 0.5 = 101 (B) 26.9 - 11.6 = 15.3 (D) 170 x 0.9 = 153 (C) 3.9 + 6.2 = 10.1 (A)